OpenStudy (anonymous):

solve. 1/3(x-2) + 1/2(x+3) = 1/4(3x+1)

OpenStudy (anonymous):

do i get common denominators first or distribute. need help thnx

OpenStudy (anonymous):

I would get common denominators first, then distribute.

OpenStudy (anonymous):

waitt i got got common denominators and then 2(6x-12)+3(3x+9) =1/4(3x=1)

OpenStudy (nikato):

\[\frac{ 1 }{ 3(x-2) }+\frac{ 1}{ 2(x+3) }= \frac{ 1 }{ 4(3x+1) }\]

OpenStudy (nikato):

i got something else. your common denominar should be 3(x-2)(X+3)(4)(3x+1)

OpenStudy (nikato):

this is so ugly. so you know x doesnt equal 2,-3, and -1/3 and you have 4(x+3)(3x+1) + 3(2)(x-2)(3x+1) = 3(x-2)(x+3) after you multiply both sides by the denominator

OpenStudy (anonymous):

oh ok

OpenStudy (nikato):

now to distribute and simpllify. do you want to try it?

OpenStudy (anonymous):

yea i will write it down and look it over

OpenStudy (anonymous):

these equations really mess w/ u

OpenStudy (nikato):

yea. they are really long. too much numbers= confusing. we could compare answers when you're done

OpenStudy (anonymous):

why was the 2 not in the common denominator with 3 and 4

OpenStudy (nikato):

because 4 is 2x2

OpenStudy (anonymous):

oh ok i see

OpenStudy (nikato):

did you copy the equation right? because i got even uglier answers

OpenStudy (anonymous):

I putting in in the common denominators

OpenStudy (anonymous):

do I distribute and then factor it

OpenStudy (nikato):

i think you should distribute and then simlify. if you factor, you would undo your distribution

OpenStudy (anonymous):

did you get the right answer

OpenStudy (nikato):

you have the answers?

OpenStudy (anonymous):

in a moment

OpenStudy (anonymous):

did you get 18^2-28x-12 after all of that distributing

OpenStudy (anonymous):

and combining

OpenStudy (nikato):

no. i got 27x^2 +7x -2 either or both of us might have made a mistake.

OpenStudy (anonymous):

I saw that you added put in a 2 with the rest of factors

OpenStudy (anonymous):

and you have 4(x+3)(3x+1) + 3(2)(x-2)(3x+1) = 3(x-2)(x+3) after you multiply both sides by the denominator

OpenStudy (anonymous):

was puttign in a (2) w/ the (x-2)(3x+1)

OpenStudy (anonymous):

a mistake

OpenStudy (nikato):

no. you need to multiply by 2. because i didnt multiply by4 to it. to get the 4 , i multiplied the 2 by 2 to get get. that why i didnt multiply by 2 with the last term

OpenStudy (nikato):